If x = cos55, y=cos 65 and z =cos 175 then xy +yz +zx =  

Grade:11

1 Answers

Soumya Narveriya
5 years ago
 
= cos55cos65+cos175cos65+cos175cos55
= cos65(cos55+cos175)+cos175cos55
= cos65(2cos115cos65)+cos175cos55   [cosA+cosB=12cos( A+B/2 )cos( A-B/2 )]
= cos65cos115+cos(180-5)cos55           [cos175=cos(180-5)=-cos5, cos60= ½ ]
=cos65cos115-cos5cos55
= ½ [2cos65cos115-2cos5cos55]
= ½ [cos180+cos50-(cos60+cos50)]      [2cosAcosB=cos(A+B)cos(A-B)]
= ½ [cos180-cos60]
= ½ [0-½]
= – ¼